# Random Access for LEO Satellite Communication Systems via Deep Learning

### v1

@@ I Introduction
Among the various types of non-terrestrial network platforms, large-scale low Earth orbit (LEO) satellite constellations have emerged as a key enabler, owing to their ability to provide near-global coverage and relatively low latency due to their proximity to the Earth compared to higher-orbit systems [WCM23_NTNMAG, WCM24_LEOhardware, Arxiv25_LEOTDD]. Reflecting this focus, since Release 17, the 3rd Generation Partnership Project (3GPP) has concentrated on adapting terrestrial cellular technologies to tackle the unique challenges of LEO satellite communication (SatCom), such as long communication distances and high mobility [36.763, 38.811, 38.821].
@@
Random access procedures are fundamental components of wireless networks that enable users to initiate connections with base stations (BSs). These procedures typically involve a sequence of message exchanges to establish uplink synchronization and request transmission resources. In terrestrial networks such as 4G LTE and 5G NR, contention-based random access is used, where users transmit randomly selected preambles and resolve potential collisions through a multi-step handshake with the BS [Book_RABO]. This approach is well-suited for terrestrial environments as it provides a scalable and efficient way to manage sporadic access attempts from a large number of users. However, adapting these contention-based mechanisms to LEO SatCom systems requires careful consideration of several LEO-specific constraints that significantly impact the performance and reliability of the access procedure. These constraints are summarized as follows, based on a representative LEO scenario with an altitude of 600 km and a minimum service elevation angle of 10 degrees [38.811]:
@@
  • Long propagation delay: The considerable distance between the satellite and ground users introduces significant propagation delays. Depending on the user’s location within the satellite footprint, the one-way delay ranges from approximately 2 ms to 6.44 ms, resulting in a round-trip time (RTT) variation of up to 4.44 ms [38.811]. In random access procedures, such delays cause latency in the exchange of signaling messages, which can degrade responsiveness and increase the time required to complete access attempts. Moreover, because the satellite BS (SBS) must receive preambles from all users without prior timing synchronization, it must set a reception window wide enough to account for this RTT uncertainty.
@@
  • Large Doppler shift: LEO satellites travel at high speeds, approximately 7.56 km/s, resulting in significant Doppler shifts. For instance, at a 2 GHz carrier frequency, this can lead to frequency offsets up to 48 kHz [38.811], which is nearly 38 times the 1.25 kHz subcarrier spacing used in the 5G NR random access procedure [38.211]. Since the random access waveform is designed under the assumption of small frequency offsets, such a large shift can cause severe inter-carrier interference and significantly degrade preamble detection performance unless compensated at the user side.
@@
  • Large number of simultaneous access attempts: Due to the large coverage area of LEO satellites compared to terrestrial BSs, the total number of users served simultaneously is expected to be significantly higher. As a result, the number of users attempting random access at the same time can be substantially greater than in terrestrial networks, leading to a higher probability of preamble collisions and access delays. For instance, in a LEO satellite-based internet of things (IoT) scenario, future systems are expected to support up to one billion IoT nodes within a single satellite footprint [36.763]. In contrast, terrestrial networks, which operate with smaller coverage areas and more densely deployed infrastructure, typically serve between 100,000 and 10,000,000 nodes in total [TWC21_DLRA, TWC17_NORA, COMST23_LEOMIMO]. This implies that satellite-based systems may need to accommodate 100 to 10,000 times more simultaneous access attempts, significantly intensifying contention on the random access channel.
@@
To address two of these challenges—namely, large RTT uncertainty and Doppler-induced frequency offset—3GPP Release 17 introduces the use of global navigation satellite system (GNSS) functionality at the user and ephemeris information broadcast by the satellite [38.811]. With knowledge of both the user’s position and the satellite’s orbit, the user can pre-compensate for the required timing advance and frequency offset before transmitting its preamble—a known sequence used to initiate random access [38.811, 38.821, 36.763]. This allows reuse of the terrestrial contention-based random access procedure in LEO SatCom systems, in which the user transmits a preamble (Step 1), the BS responds (Step 2), the user sends identity information (Step 3), and the BS acknowledges (Step 4). However, despite enabling this reuse, the approach suffers from critical limitations in detecting and resolving preamble collisions. Collisions in Step 1 are not detected at the time of transmission but are only revealed later when the BS fails to decode signals in Step 3. On the user side, a collision is inferred only when a message is not received within the collision resolution (CR) timer in Step 4.
@@
This mechanism results in delays of at least two RTTs and leads to inefficient use of physical uplink shared channel (PUSCH) resources allocated for Step 3. Such issues are present in terrestrial networks but exacerbate in LEO systems due to inherently longer delays and a larger number of simultaneous access attempts. LEO satellite-based IoT scenarios currently serve as a representative example, where the high density of devices leads to frequent access contention. Similar challenges are also expected in emerging applications such as direct-to-device (D2D) services, where the number of connected handheld devices is projected to grow significantly in the coming years [AESM25_D2D]. These observations highlight the need for improved random access mechanisms that can detect collisions early and increase access success rates, particularly in LEO SatCom systems operating under high user density.
@@ II-A1 Reducing the likelihood of preamble collisions
To alleviate preamble collisions in contention-based access, various techniques have been proposed to reduce the probability that multiple users select the same preamble simultaneously. Key approaches explored in the literature include access class barring (ACB) [TVT18_ACB1, TVT23_ACB2, TWC24_ACB3], adaptive random backoff strategies [TVT16_BO1, TN18_BO2, TMC24_BO3], and the use of non-orthogonal preambles [TVT18_NP1, TVT21_NP2, TWC21_NP3, CL17_NP4, TVT23_TwoStage] have received particular attention.
In the ACB scheme, the number of access attempts per random access channel (RACH) slot is regulated to avoid excessive collisions. Before transmitting a preamble, each user performs an ACB check and proceeds only if it passes; otherwise, the user defers the attempt to the next available RACH slot [TVT18_ACB1, TVT23_ACB2, TWC24_ACB3]. Similarly, many works have focused on adjusting the backoff behavior of users after collisions. In [TVT16_BO1, TN18_BO2, TMC24_BO3], dynamic backoff schemes were proposed that adapt the uniform backoff window based on observed collision levels. While ACB and backoff adjustment help mitigate RACH congestion, they require tracking the backlog state of all users. These approaches are less attractive in LEO SatCom systems, where the service region changes dynamically due to satellite movement.
One of the most effective ways to reduce preamble collisions is to increase the number of available preambles [TVT18_NP1, TVT21_NP2, TWC21_NP3, CL17_NP4, TVT23_TwoStage]. The authors of [TVT18_NP1, TVT21_NP2] proposed a concatenated preamble structure that forms a composite sequence by combining two Zadoff-Chu (ZC) sequences generated from different roots. This combinatorial design increases the number of unique preambles without requiring longer sequences or additional time-frequency resources. Although this expanded preamble space reduces collision probability, it also leads to increased non-orthogonality, resulting in higher interference as the user density grows. To mitigate such interference, successive interference cancellation techniques have been employed in [TWC21_NP3, CL17_NP4, TVT23_TwoStage]. While these methods help detect non-orthogonal preambles more effectively, they rely on iterative cancellation procedures, which entail high computational complexity.
@@ II-A2 Resolving preamble collisions once they occur
When multiple users select the same preamble and transmit it simultaneously, the BS cannot distinguish between them and responds with the same message. As a result, all those users receive the same grant for Step 3, causing a collision that leads to random access failure. Several approaches have been proposed to handle such collisions after they occur, aiming to recover from contention rather than prevent it in advance. These include transmission control [PIMRC17_BCCR, IoTJ20_MSG3B], non-orthogonal random access [TCOM23_NOMA1, IoTJ23_NOMA2, TMC24_NOMA3], and early preamble collision detection schemes [TWC17_NORA, TVT18_eCD2, TVT21_SeCD, TVT25_MTeCD, TWC21_DLRA, TWC23_DLDouble].
@@
A user-side approach in [PIMRC17_BCCR] enables early collision detection by having users broadcast short contention signals before Step 3 transmission. While effective in terrestrial settings, this method is unsuitable for LEO SatCom due to wide coverage and limited sensing among users. Alternatively, a BS-centric scheme called Step 3 barring was proposed in [IoTJ20_MSG3B], where the BS estimates collision levels and broadcasts a uniform barring probability to all users. Although compatible with LEO SatCom, this approach cannot differentiate collided from non-collided users, causing unnecessary barring and increased access delay.
@@
In [TCOM23_NOMA1, IoTJ23_NOMA2, TMC24_NOMA3], non-orthogonal random access schemes were proposed where users sharing the same Step 3 grant transmit simultaneously using power-domain NOMA. This enables multiple users to share uplink resources through superimposed transmissions and works well in terrestrial networks with diverse channel gains. In LEO SatCom, however, the long transmission distance yields minimal power disparities, making such separation challenging and limiting their effectiveness [ICASSP23_NOMA].
@@
Several studies have sought to detect preamble collisions at the BS immediately after receiving Step 1 transmissions. A common approach is to analyze the correlation output to identify multiple time of arrival (ToA) peaks [TWC17_NORA, TVT18_eCD2], which works in terrestrial networks but becomes ineffective in LEO SatCom due to timing advance pre-compensation using GNSS and ephemeris information [36.763]. To overcome this limitation, alternative schemes that do not rely on ToA differences have been proposed [TVT21_SeCD, TVT25_MTeCD, TWC21_DLRA, TWC23_DLDouble]. For example, a prime-based cyclic shift scheme [TVT21_SeCD] enables fast collision detection but is sensitive to RTT uncertainty and fading. A multi-threshold method [TVT25_MTeCD] instead uses empirically derived thresholds, but lacks robustness under varying channel conditions.
@@
There have also been attempts to apply deep learning for early preamble collision detection [TWC21_DLRA, TWC23_DLDouble]. In [TWC21_DLRA], a neural network was used to estimate the number of colliding users and extract timing advance values, enabling the BS to broadcast multiple Step 3 grants. However, in LEO SatCom, users perform timing advance pre-compensation and the satellite is moving, so fixed timing advance values are unavailable, making this approach impractical. Although the method shows strong detection performance, its fully connected network incurs high complexity. In [TWC23_DLDouble], a double contention scheme was proposed in which a neural network allocates two Step 3 grants when two users collide, but it also suffers from high complexity, limited scalability beyond two users, and the need for dynamically adjustable Step 3 resources, which conflicts with current standards [Book_RABO].
@@ II-B Contributions
  • We propose a deep learning-based early preamble collision classifier tailored for LEO SatCom systems. The classifier leverages correlation values from multiple antennas or ports as input features, enabling early estimation of the number of collided users during the initial stage of the random access process. To reduce model complexity while maintaining classification accuracy, a lightweight 1D convolutional neural network architecture is adopted. The key advantage of our approach lies in its ability to enable fast contention resolution with a lightweight model, making it well-suited for practical deployment in LEO SatCom systems.
  • We propose an opportunistic transmission scheme for Step 3 to maximize random access success probability. The scheme exploits the estimated number of collided users provided by the early preamble collision classifier to compute an access probability for each preamble index. This access probability is optimized to balance the trade-off between reducing unnecessary Step 3 transmissions and ensuring successful access opportunities for collided users. By allowing users to attempt Step 3 opportunistically rather than enforcing a strict backoff, the scheme enhances random access success probability, minimizes access delay, and preserves uplink resource utilization.
  • We provide simulation results to demonstrate the effectiveness of the proposed framework. The simulations are conducted under parameter settings reflecting representative LEO SatCom scenarios defined by 3GPP. The results show that the proposed framework outperforms conventional approaches and prior works in terms of classification accuracy, computational complexity, PUSCH utilization, the average random access delay, and the number of successful random access users.
@@ III Preliminaries
This work builds upon the contention-based random access framework defined by 3GPP for SatCom systems [38.811]. Before introducing our proposed framework, it is therefore useful to outline the 3GPP-specified procedure to establish notation and terminology. As illustrated in Fig. 1, the contention-based random access framework for LEO SatCom systems consists of the following four steps.
@@
(Step 1) Preamble Transmission: Since 4G LTE systems, a ZC sequence has been used as the random access preamble in 3GPP [Book_RABO], and is expressed as
@@
The ZC sequence also has the following cyclic auto-correlation property in which the auto-correlation function takes the form of the Dirac delta function $\delta[\cdot]$ as
@@
where $N_{\mathrm{CS}}$ represents the cyclic shift size and determines the range of $i\in\mathopen{}\mathclose{{\left\{0,...,\mathopen{}\mathclose{{\left(\lfloor N_{\mathrm{ZC}}/N_{\mathrm{CS}}\rfloor-1}}\right)}}\right\}$. Since there are $N_{\mathrm{R}}$ root indices, the total number of available preambles is $N_{\mathrm{PA}}=N_{\mathrm{R}}\cdot\lfloor N_{\mathrm{ZC}}/N_{\mathrm{CS}}\rfloor$. Users randomly select one of these available preambles and transmit it to the SBS when requesting random access. Given the large RTT uncertainty and Doppler shift inherent in satellite links, users are assumed to apply pre-compensation for timing advance and frequency offset prior to transmission [38.811]. As investigated in 3GPP, such compensation can be accomplished by GNSS-based synchronization, which can achieve high estimation accuracy [38.133, 3GPP_FreqEst].
@@
(Step 2) Preamble Detection and Random Access Response Transmission: After users transmit their preambles, the SBS receives the composite signal and computes the correlation values between the received signal and each of the $N_{\mathrm{R}}$ root ZC sequences to identify the transmitted preambles, as illustrated in Fig. 2. Each root sequence is associated with multiple cyclically shifted versions, and each of these shifted sequences defines a preamble. To enable accurate detection, a ZCZ is configured for each root sequence, within which the autocorrelation of the ZC sequence is ideally zero, minimizing mutual interference among different preambles. The correlation value is computed within this ZCZ region, and if the maximum value exceeds a predefined threshold, the corresponding preamble is considered to be detected. In this work, the ZCZ length is set equal to the cyclic shift size $N_{\mathrm{CS}}$, ensuring that each user’s preamble falls within a distinct non-overlapping zone and can be reliably detected by the SBS. In terrestrial networks, the ZCZ length is typically determined based on the maximum expected timing misalignment, which includes the maximum timing advance and the channel delay spread. However, in LEO SatCom systems, users perform timing advance pre-compensation before transmission, which alters the delay characteristics. As a result, the design of ZCZ in SatCom must consider this pre-compensation behavior, and the required ZCZ length is derived accordingly.
@@
(Step 4) Contention Resolution Message Transmission: If the SBS successfully decodes the RRC connection request message, it transmits a CR message containing the user identifier (ID). If multiple users simultaneously use the same PUSCH resource, the SBS would presumably fail to decode the RRC connection request messages in that resource and would thus not respond to those requests. A user that receives the CR message within the predefined CR timer sends back a positive acknowledgement (ACK). In contrast, users that do not receive the CR message before the CR timer expires reinitiate the random access procedure after some delay based on a predefined random backoff policy [Book_RABO].

### v2

@@ I Introduction
Among the various types of non-terrestrial network platforms, large-scale low Earth orbit (LEO) satellite constellations have emerged as a key enabler, owing to their ability to provide near-global coverage and relatively low latency due to their proximity to the Earth compared to higher-orbit systems [17, 41, 28]. Reflecting this focus, since Release 17, the 3rd Generation Partnership Project (3GPP) has concentrated on adapting terrestrial cellular technologies to tackle the unique challenges of LEO satellite communication (SatCom), such as long communication distances and high mobility [6, 7, 5].
@@
Random access procedures are fundamental components of wireless networks that enable users to initiate connections with base stations (BSs). These procedures typically involve a sequence of message exchanges to establish uplink synchronization and request transmission resources. In terrestrial networks such as 4G LTE and 5G NR, contention-based random access is used, where users transmit randomly selected preambles and resolve potential collisions through a multi-step handshake with the BS [14]. This approach is well-suited for terrestrial environments as it provides a scalable and efficient way to manage sporadic access attempts from a large number of users. However, adapting these contention-based mechanisms to LEO SatCom systems requires careful consideration of several LEO-specific constraints that significantly impact the performance and reliability of the access procedure. These constraints are summarized as follows, based on a representative LEO scenario with an altitude of 600 km and a minimum service elevation angle of 10 degrees [7]:
@@
  • Long propagation delay: The considerable distance between the satellite and ground users introduces significant propagation delays. Depending on the user’s location within the satellite footprint, the one-way delay ranges from approximately 2 ms to 6.44 ms, resulting in a round-trip time (RTT) variation of up to 4.44 ms [7]. In random access procedures, such delays cause latency in the exchange of signaling messages, which can degrade responsiveness and increase the time required to complete access attempts. Moreover, because the satellite BS (SBS) must receive preambles from all users without prior timing synchronization, it must set a reception window wide enough to account for this RTT uncertainty.
@@
  • Large Doppler shift: LEO satellites travel at high speeds, approximately 7.56 km/s, resulting in significant Doppler shifts. For instance, at a 2 GHz carrier frequency, this can lead to frequency offsets up to 48 kHz [7], which is nearly 38 times the 1.25 kHz subcarrier spacing used in the 5G NR random access procedure [9]. Since the random access waveform is designed under the assumption of small frequency offsets, such a large shift can cause severe inter-carrier interference and significantly degrade preamble detection performance unless compensated at the user side.
@@
  • Large number of simultaneous access attempts: Due to the large coverage area of LEO satellites compared to terrestrial BSs, the total number of users served simultaneously is expected to be significantly higher. As a result, the number of users attempting random access at the same time can be substantially greater than in terrestrial networks, leading to a higher probability of preamble collisions and access delays. For instance, in a LEO satellite-based internet of things (IoT) scenario, future systems are expected to support up to one billion IoT nodes within a single satellite footprint [6]. In contrast, terrestrial networks, which operate with smaller coverage areas and more densely deployed infrastructure, typically serve between 100,000 and 10,000,000 nodes in total [20, 30, 18]. This implies that satellite-based systems may need to accommodate 100 to 10,000 times more simultaneous access attempts, significantly intensifying contention on the random access channel. Notably, even when a phased-array antenna confines contention to a narrow beam, the beam footprint remains much larger than that of a typical terrestrial cell [18], and thus high contention can still arise within a single beam.
@@
To address two of these challenges—namely, large RTT uncertainty and Doppler-induced frequency offset—3GPP Release 17 introduces the use of global navigation satellite system (GNSS) functionality at the user and ephemeris information broadcast by the satellite [7]. With knowledge of both the user’s position and the satellite’s orbit, the user can pre-compensate for the required timing advance and frequency offset before transmitting its preamble—a known sequence used to initiate random access [7, 5, 6]. This allows reuse of the terrestrial contention-based random access procedure in LEO SatCom systems, in which the user transmits a preamble (Step 1), the BS responds (Step 2), the user sends identity information (Step 3), and the BS acknowledges (Step 4). However, despite enabling this reuse, the approach suffers from critical limitations in detecting and resolving preamble collisions. Collisions in Step 1 are not detected at the time of transmission but are only revealed later when the BS fails to decode signals in Step 3. On the user side, a collision is inferred only when a message is not received within the collision resolution (CR) timer in Step 4.
@@
This mechanism results in delays of at least two RTTs and leads to inefficient use of physical uplink shared channel (PUSCH) resources allocated for Step 3. Such issues are present in terrestrial networks but exacerbate in LEO systems due to inherently longer delays and a larger number of simultaneous access attempts. LEO satellite-based IoT scenarios currently serve as a representative example, where the high density of devices leads to frequent access contention. Similar challenges are also expected in emerging applications such as direct-to-device (D2D) services, where the number of connected handheld devices is projected to grow significantly in the coming years [13]. These observations highlight the need for improved random access mechanisms that can detect collisions early and increase access success rates, particularly in LEO SatCom systems operating under high user density.
@@ II-A1 Reducing the likelihood of preamble collisions
To alleviate preamble collisions in contention-based access, various techniques have been proposed to reduce the probability that multiple users select the same preamble simultaneously. Key approaches explored in the literature include access class barring (ACB) [39, 32, 15], adaptive random backoff strategies [31, 12, 33], and non-orthogonal preambles [19, 34, 24, 25, 21].
In the ACB scheme, the number of access attempts per random access channel (RACH) slot is regulated to avoid excessive collisions. Before transmitting a preamble, each user performs an ACB check and proceeds only if it passes; otherwise, the user defers the attempt to the next available RACH slot [39, 32, 15]. Similarly, many works have focused on adjusting the backoff behavior of users after collisions. In [31, 12, 33], dynamic backoff schemes were proposed that adapt the uniform backoff window based on observed collision levels. While ACB and backoff adjustment help mitigate RACH congestion, they require tracking the backlog state of all users. These approaches are less attractive in LEO SatCom systems, where the service region changes dynamically due to satellite movement.
One of the most effective ways to reduce preamble collisions is to increase the number of available preambles [19, 34, 24, 25, 21]. The authors of [19, 34] proposed a concatenated preamble structure that forms a composite sequence by combining two Zadoff-Chu (ZC) sequences generated from different roots. This combinatorial design increases the number of unique preambles without requiring longer sequences or additional time-frequency resources. Although this expanded preamble space reduces collision probability, it also leads to increased non-orthogonality, resulting in higher interference as the user density grows. To mitigate such interference, successive interference cancellation techniques have been employed in [24, 25, 21]. While these methods help detect non-orthogonal preambles more effectively, they rely on iterative cancellation procedures, which entail high computational complexity.
Recent learning-based studies have investigated access-control policies for LEO satellite networks to mitigate contention under non-stationary topology and time-varying multi-satellite coverage. For example, [29] proposes an emergent contention-based random access protocol (eRACH) learned via multi-agent DRL, where distributed agents adapt their access behaviors to evolving network states. Similarly, [42] formulates user access as a long-term decision problem under time-varying coverage and develops a decentralized multi-agent RL mechanism to maximize long-term throughput. In addition, some DRL studies in the satellite context focus on higher-layer network control such as transport-layer congestion control in space–air–ground integrated networks (SAGINs), rather than contention-based random access [16]. Overall, these RL-based approaches aim to reduce congestion through long-term access adaptation, thereby lowering the likelihood of persistent contention in random access procedures.
@@ II-A2 Resolving preamble collisions once they occur
When multiple users select the same preamble and transmit it simultaneously, the BS cannot distinguish between them and responds with the same message. As a result, all those users receive the same grant for Step 3, causing a collision that leads to random access failure. Several approaches have been proposed to handle such collisions after they occur, aiming to recover from contention rather than prevent it in advance. These include transmission control [40, 35], non-orthogonal random access [37, 36, 23], and early preamble collision detection schemes [30, 11, 44, 22, 20, 43].
@@
A user-side approach in [40] enables early collision detection by having users broadcast short contention signals before Step 3 transmission. While effective in terrestrial settings, this method is unsuitable for LEO SatCom due to wide coverage and limited sensing among users. Alternatively, a BS-centric scheme called Step 3 barring was proposed in [35], where the BS estimates collision levels and broadcasts a uniform barring probability to all users. Although compatible with LEO SatCom, this approach cannot differentiate collided from non-collided users, causing unnecessary barring and increased access delay.
@@
In [37, 36, 23], non-orthogonal random access schemes were proposed where users sharing the same Step 3 grant transmit simultaneously using power-domain NOMA. This enables multiple users to share uplink resources through superimposed transmissions and works well in terrestrial networks with diverse channel gains. In LEO SatCom, however, the long transmission distance yields minimal power disparities, making such separation challenging and limiting their effectiveness [38].
@@
Several studies have sought to detect preamble collisions at the BS immediately after receiving Step 1 transmissions. A common approach is to analyze the correlation output to identify multiple time of arrival (ToA) peaks [30, 11], which works in terrestrial networks but becomes ineffective in LEO SatCom due to timing advance pre-compensation using GNSS and ephemeris information [6]. To overcome this limitation, alternative schemes that do not rely on ToA differences have been proposed [44, 22, 20, 43]. For example, a prime-based cyclic shift scheme [44] enables fast collision detection but is sensitive to RTT uncertainty and fading. A multi-threshold method [22] instead uses empirically derived thresholds, but lacks robustness under varying channel conditions.
@@
There have also been attempts to apply deep learning for early preamble collision detection [20, 43]. In [20], a neural network was used to estimate the number of colliding users and extract timing advance values, enabling the BS to broadcast multiple Step 3 grants. However, in LEO SatCom, users perform timing advance pre-compensation and the satellite is moving, so fixed timing advance values are unavailable, making this approach impractical. Although the method shows strong detection performance, its fully connected network incurs high complexity. In [43], a double contention scheme was proposed in which a neural network allocates two Step 3 grants when two users collide, but it also suffers from high complexity, limited scalability beyond two users, and the need for dynamically adjustable Step 3 resources, which conflicts with current standards [14].
@@ II-B Contributions
  • We propose a deep learning-based early preamble collision classifier tailored for LEO SatCom systems. The classifier leverages correlation values from multiple antennas or ports as input features, enabling early estimation of the collision multiplicity (i.e., the number of users selecting the same preamble) during the initial stage of the random access process. To satisfy practical satellite payload constraints, a lightweight 1D convolutional neural network architecture is adopted to reduce on-board computation and memory requirements while maintaining high classification accuracy. This lightweight design enables fast contention handling and supports practical NTN deployment under stringent power and latency budgets.
  • We propose a collision-aware opportunistic Step 3 transmission mechanism to improve random access performance. The scheme exploits the estimated collision multiplicity from the proposed classifier and computes a preamble-index-wise access probability that explicitly controls the number of Step 3 attempts. The access probability is designed to balance the trade-off between suppressing redundant Step 3 transmissions and preserving successful access opportunities for collided users. By enabling probabilistic Step 3 attempts rather than a strict backoff, the proposed mechanism improves the random access success opportunity and mitigates delay and uplink resource waste caused by unnecessary transmissions.
  • We provide an NTN-compliant, system-level validation under a 3GPP-aligned LEO SatCom scenario and extract actionable design insights. Rather than merely presenting simulation plots, the evaluation substantiates the feasibility of integrating the proposed framework into a 3GPP-aligned random access procedure and quantifies its system-level gains over benchmark schemes. In addition, the results identify operating regimes where the proposed framework yields clear benefits and offer practical guidelines for key system parameters such as active-user load and received SNR regimes.
@@ III Preliminaries
This work builds upon the contention-based random access framework defined by 3GPP for SatCom systems [7]. Before introducing our proposed framework, it is therefore useful to outline the 3GPP-specified procedure to establish notation and terminology. As illustrated in Fig. 1, the contention-based random access framework for LEO SatCom systems consists of the following four steps.
@@
(Step 1) Preamble Transmission: Since 4G LTE systems, a ZC sequence has been used as the random access preamble in 3GPP [14], and is expressed as
@@
Here, $c_{rs}[m]$ denotes the normalized cyclic cross-correlation, where the unnormalized correlation $\sum_{n=0}^{N_{\mathrm{ZC}}-1}z_{r}[n]z_{s}^{*}[(n+m)\bmod N_{\mathrm{ZC}}]$ has magnitude $\sqrt{N_{\mathrm{ZC}}}$. The ZC sequence also has the following cyclic auto-correlation property in which the auto-correlation function takes the form of the Dirac delta function $\delta[\cdot]$ as
@@
where $N_{\mathrm{CS}}$ represents the cyclic shift size and determines the range of $i\in\mathopen{}\mathclose{{\left\{0,...,\mathopen{}\mathclose{{\left(\lfloor N_{\mathrm{ZC}}/N_{\mathrm{CS}}\rfloor-1}}\right)}}\right\}$. Since there are $N_{\mathrm{R}}$ root indices, the total number of available preambles is $N_{\mathrm{PA}}=N_{\mathrm{R}}\cdot\lfloor N_{\mathrm{ZC}}/N_{\mathrm{CS}}\rfloor$. Users randomly select one of these available preambles and transmit it to the SBS when requesting random access. Given the large RTT uncertainty and Doppler shift inherent in satellite links, users are assumed to apply pre-compensation for timing advance and frequency offset prior to transmission [7]. As investigated in 3GPP, such compensation can be accomplished by GNSS-based synchronization, which can achieve high estimation accuracy [10, 4].
@@
(Step 2) Preamble Detection and Random Access Response Transmission: After users transmit their preambles, the SBS receives the composite signal and computes the correlation values between the received signal and each of the $N_{\mathrm{R}}$ root ZC sequences to identify the transmitted preambles, as illustrated in Fig. 2. Each root sequence is associated with multiple cyclically shifted versions, and each of these shifted sequences defines a preamble. To enable accurate detection, a zero correlation zone (ZCZ) is configured for each root sequence, within which the autocorrelation of the ZC sequence is ideally zero, minimizing mutual interference among different preambles. The correlation value is computed within this ZCZ region, and if the maximum value exceeds a predefined threshold, the corresponding preamble is considered to be detected. In this work, the ZCZ length is set equal to the cyclic shift size $N_{\mathrm{CS}}$, ensuring that each user’s preamble falls within a distinct non-overlapping zone and can be reliably detected by the SBS. In terrestrial networks, the ZCZ length is typically determined based on the maximum expected timing misalignment, which includes the maximum timing advance and the channel delay spread. However, in LEO SatCom systems, users perform timing advance pre-compensation before transmission, which alters the delay characteristics. As a result, the design of ZCZ in SatCom must consider this pre-compensation behavior, and the required ZCZ length is derived accordingly.
@@
(Step 4) Contention Resolution Message Transmission: If the SBS successfully decodes the RRC connection request message, it transmits a CR message containing the user identifier (ID). If multiple users simultaneously use the same PUSCH resource, the SBS would presumably fail to decode the RRC connection request messages in that resource and would thus not respond to those requests. A user that receives the CR message within the predefined CR timer sends back a positive acknowledgement (ACK). In contrast, users that do not receive the CR message before the CR timer expires reinitiate the random access procedure after some delay based on a predefined random backoff policy [14].